OpenStudy (anonymous):

I'm confused about what this question is asking me to do: In exercise 17, use the Lagrange form of the remainder to prove that the Maclaurin series converges to the generating function from the given exercise (#7). #7 just gives you xe^x, so I don't understand what I am supposed to find the error with

OpenStudy (amistre64):

find the general error for the n+1th term i spose. can you provide a pic of the question?

OpenStudy (amistre64):

what is the maclaruin series for x e^x?

OpenStudy (anonymous):

x^n+1/n!

OpenStudy (amistre64):

\[e^x=\sum_{n=0}^{\infty}\frac{1}{n!}x^n\] \[xe^x=\sum_{n=0}^{\infty}\frac{1}{n!}x^{n+1}\]

OpenStudy (amistre64):

how do we determine the formula for truncation error with this now?

OpenStudy (anonymous):

Substitute n+1 in for n

OpenStudy (amistre64):

right, and?

OpenStudy (anonymous):

Take the derivative at the maximum value on the interval?

OpenStudy (anonymous):

Except for the fact that I was not provided an interval!

OpenStudy (amistre64):

well, its valid for all x in R isnt it?

OpenStudy (anonymous):

Yes, but you have to take the derivative at c, which is the highest point on the interval

OpenStudy (amistre64):

well, for a given interval [a,b] the highest point is b

OpenStudy (anonymous):

No, its whatever makes the nth derivative largest

OpenStudy (amistre64):

well, prolly not accurate, but its a monotonic function for x=0 and above ...

OpenStudy (amistre64):

f = x e^x f' = x' e^x + x e^x = (1+x) e^x f'' = (1+x)' e^x + (1+x) e^x = (2+x) e^x f''' = (3+x) e^x f^(n) = (n+x)e^x right?

OpenStudy (amistre64):

how do we determine what makes the nth derivative 'largest' with this?

OpenStudy (anonymous):

I guess it is infinity

OpenStudy (amistre64):

e^x dominates the linear factor .... im guess e^x is max as x to infinity? then we need to prolly limit if c to infinity .. but thats just a guess.

OpenStudy (anonymous):

I have absolutely no idea. My book is absolutely terrible. I had no idea what the question was asking in the first place

OpenStudy (amistre64):

seems to be something like this \[\lim_{c\to\infty}R_n=\lim_{c\to\infty}\left|\frac{(c+n)e^c~c^{n+1}}{(n+1)!}\right|\]

OpenStudy (amistre64):

not = but <=

OpenStudy (anonymous):

We never learned anything like that

OpenStudy (amistre64):

neither have i, but then i havent taken an error estimation course or a course with it in it.

OpenStudy (anonymous):

This is calc 2

OpenStudy (anonymous):

Or BC calc, whichever you prefer

OpenStudy (amistre64):

https://www.youtube.com/watch?v=cMc6l8AhF9M

OpenStudy (anonymous):

I know what Lagrange Error is; my confusion is how I am supposed to do it without an interval. I have no idea what they are referring to when they say the Maclaurin series or the generating function. They only gave me one equation

OpenStudy (amistre64):

the link i sent explains it, im watching it now .. its a 10-15min video

OpenStudy (anonymous):

I finished the video

OpenStudy (amistre64):

we hav a Mac series, so its centered at 0 \[xe^x=\sum_0^{k}\frac{x^{n+1}}{n!}+R_n\] \[f^{(n)}[xe^x]=(n+x)e^x\implies R_n\le\frac{(n+1+c)e^cx^{n+2}}{(n+1)!}\] \[\lim_{n\to \infty} \frac{(n+1+c)e^c x^{n+2}}{(n+1)!}\] \[e^c~\lim_{n\to \infty} \frac{(n+1+c)x^{n+2}}{(n+1)!}\] if the limit goes to 0, then the Remainder converges assuming my f^(n)(c) is correct

OpenStudy (anonymous):

Oooooh, I see. So I guess that they want me to prove that xe^x converges to the series that I wrote for xe^x?

OpenStudy (amistre64):

prove that the limit of the Remainder goes to 0 as n approaches infinity yes

OpenStudy (anonymous):

OOOH, I GET WHAT THE QUESTION IS ASKING NOW!

OpenStudy (anonymous):

Thanks so much for helping me!!!

OpenStudy (amistre64):

youre welcome, sorry it took so long, but i had to basically learn it along the way :)

OpenStudy (anonymous):

That was a horribly worded question. I really appreciate it